ercise 9-2 (Algorithmic) (LO. 2) Compute the standard deduction for the following taxpayers. If an amount...

  1. ercise 9-2 (Algorithmic) (LO. 2)

    Compute the standard deduction for the following taxpayers.

    If an amount is zero, enter "0".

    Click here to access the standard deduction table to use.

    a. Margie is 15 and claimed as a dependent by her parents. She has $475 in dividend income and $2,520 in wages from a part-time job. $
    b. Ruby and Woody are married and file a joint tax return. Ruby is age 66, and Woody is 69. Their taxable retirement income is $11,260. $
    c. Shonda is age 68 and single. She is claimed by her daughter as a dependent. Her earned income is $180, and her interest income is $605. $
    d. Frazier, age 50, is married but is filing a separate return. His wife itemizes her deductions. $

Check My Work5 more Check My Work uses remaining.

  • Next
Answer #1

Part A

Standard deduction greater of 1050 or (earned income +350)

Therefore,

Standard deduction = 2520+350 = 2870

Part B

Their total standard deduction = 24000+2600 (As they both are over 65 years) = 26600

Part C

For dependent standard deduction is greater of 1050 or (earned income +350), if filing independently

When claimed as dependent, it will change. It will be

Therefore,

Standard deduction = 1050+1600 = 2650

Part D

Standard deduction for married filing separately = 0 (because he is not eligible for standard deduction and thus must itemize as his wife itemizes)